Find a basis for the orthognonal complement of U=span Lona i Number of Vectors: 1 Basis for Uł:
Answer #1

Griveno u = spama 3 we have to find basis for the orthogonal complement of u. i. e. for ut ut = { x = (a, b, c, d) EIRY : x =5atabt act 2d = 0 al- o -3d=0 > BIS d=0 -2b - 20 C o as d=0 reduces to 5a + 2b +2c=0 - a=-26 - 20 = 65-215 7 cc-2157 - SO dho